OpenStudy (anonymous):

What is the simplified form of the cube root of x squared times the third root of x squared

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

\[(\sqrt[3]{x})^{2} \times (\sqrt[3]{x})^{2} \]

OpenStudy (nikato):

is it this?

OpenStudy (anonymous):

@nikato Yes But The 2 Is A Exponent For The X Under Tha Square Root Symbol

OpenStudy (nikato):

for both of them?

OpenStudy (anonymous):

Yes

OpenStudy (nikato):

\[(\sqrt[3]{x ^{2}}) \times (\sqrt[3]{x ^{2}})\]

OpenStudy (anonymous):

Yes

OpenStudy (nikato):

well you can rewrite the square roots to exponents

OpenStudy (nikato):

\[x ^{\frac{ 2 }{ 3 }}\times x ^{\frac{ 2 }{ 3 }}\]

OpenStudy (nikato):

so  and simplify

OpenStudy (anonymous):

Would It Be... x^2/3

OpenStudy (anonymous):

Because 2+2=4 And 3+3=6 So 4/2=2 2/6=3

OpenStudy (nikato):

remember you dont add the denominator so x^4/3

OpenStudy (anonymous):

Isnt It X^4/9

OpenStudy (nikato):

so i guess you can leave it like this or put it back with square roots